# Learning to accelerate distributed ADMM using graph neural networks

### v1

@@ D.2 Message-passing formulation of distributed ADMM
$2\sum_{j\in\mathcal{N}(i)\cup\{i\}}(P_{ji}(x_{i}-x_{i}^{k}))^{T}y_{j}^{k}=2(x_{i}-x_{i}^{k})^{T}\Biggl{(}P_{ii}^{T}y_{i}^{k}+\underbrace{\sum_{j\in\mathcal{N}(i)}P_{ji}^{T}y_{j}^{k}}_{=:\,\bar{y}_{\rightarrow i}^{\,k}}\Biggl{)}\,$, (18)
@@
2:  $m_{ji}\leftarrow\left(P_{ji}\lambda_{j}^{k},P_{ji}y_{j}^{k}\right)$ $\triangleright$ message 1
@@
5:  $\left(\bar{\lambda}_{\rightarrow i}^{k},\bar{y}_{\rightarrow i}^{k}\right)\leftarrow\sum_{(j,i)\in\mathcal{E}}m_{ji}$ $\triangleright$ aggregation 1
6:     ${v}_{i}\hskip 19.34766pt\leftarrow\left(x_{i}^{k+1},y_{i}^{k},\lambda_{i}^{k}\right),\,\text{where }\triangleright$ update 1
@@
12:  $\bar{x}_{\rightarrow i}^{k+1}\leftarrow\sum_{(j,i)\in\mathcal{E}}m_{ji}$ $\triangleright$ aggregation 2
13:   ${v}_{i}\hskip 6.544pt\leftarrow\left(x_{i}^{k+1},y_{i}^{k+1},\lambda_{i}^{k+1}\right),\,\text{where }$ $\triangleright$ update 2

### v2

@@ E.2 Message-passing formulation of distributed ADMM
$2\cdot\sum_{j\in\mathcal{N}(i)\cup\{i\}}(P_{ji}(x_{i}-x_{i}^{k}))^{T}y_{j}^{k}=2(x_{i}-x_{i}^{k})^{T}\Biggl(P_{ii}^{T}y_{i}^{k}+\underbrace{\sum_{j\in\mathcal{N}(i)}P_{ji}^{T}y_{j}^{k}}_{=:\,\bar{y}_{\rightarrow i}^{\,k}}\Biggl)\,$, (17)
@@
2:  $m_{ji}=\left(P_{ji}\lambda_{j}^{k},P_{ji}y_{j}^{k}\right)$ $\triangleright$ message 1
@@
5:  $\left(\bar{\lambda}_{\rightarrow i}^{k},\bar{y}_{\rightarrow i}^{k}\right)=\sum_{(j,i)\in\mathcal{E}}m_{ji}$ $\triangleright$ aggregation 1
6:     ${v}_{i}\hskip 22.47748pt=\left(x_{i}^{k+1},y_{i}^{k},\lambda_{i}^{k}\right),\,\text{where }\triangleright$ update 1
@@
12:  $\bar{x}_{\rightarrow i}^{k+1}=\sum_{(j,i)\in\mathcal{E}}m_{ji}$ $\triangleright$ aggregation 2
13:   ${v}_{i}\hskip 7.96674pt=\left(x_{i}^{k+1},y_{i}^{k+1},\lambda_{i}^{k+1}\right),\,\text{where }$ $\triangleright$ update 2